OpenStudy (sydthekid913):

What is the first step of equations using elimination? y=3/4x+3; 2x-4y=6

OpenStudy (sydthekid913):

Will medal best answer

OpenStudy (sydthekid913):

A. Write the second equation in slope-intercept form B. Subtract 2x from both sides of the second equation C. Write the first equation in standard form D. Multiply the first equation by 3

OpenStudy (texaschic101):

y = 3/4x + 3 -- change to standard form -3/4x + y = 3 -- multiply by -1 to make x positive 3/4x - y = -3

OpenStudy (texaschic101):

answer C

OpenStudy (sydthekid913):

Thank you but why -1?

OpenStudy (texaschic101):

slope intercept form is y = mx + b...not -y - mx + b. -3/4x + y = 3 --- multiply by -1 (-1) -3/4x + (-1)y = (-1)3 3/4x - y = -3

OpenStudy (texaschic101):

I suppose you could still do the problem even if you left x negative

OpenStudy (sydthekid913):

So how would I solve a problem like this 2x+3y=13; -6x+4y=0?

OpenStudy (texaschic101):

ooops..typo up there I meant to say standard form is Ax + By = C and x needs to be positive. ======================= 2x + 3y = 13 --->(3) -6x + 4y = 0 -------------- 6x + 9y = 39 (result of multiplying by 3) -6x + 4y = 0 --------------add 13y = 39 y = 39/13 y = 3 now sub 3 in for y in either equation to find x 2x + 3y = 13 2x + 3(3) = 13 2x + 9 = 13 2x = 13 - 9 2x = 4 x = 2 check.. -6x + 4y = 0 -6(2) + 4(3) = 0 -12 + 12 = 0 0 = 0 (correct) so x = 2 and y = 3

OpenStudy (sydthekid913):

Wow thank you so much!

OpenStudy (texaschic101):

no problem :)

OpenStudy (sydthekid913):

Hahaha just lost my tooth